Two cars are travelling in the same direction. Car A is 532m in front of Car B, is travelling at 54km/h. Car B is travelling at 60km/h. Calculate how long it takes for Car B to catch up to Car A.
1
Expert's answer
2019-01-15T09:05:11-0500

d_1=v_1 t+d

d_2=v_2 t

Thus,

d_1=d_2

v_1 t+d=v_2 t

54/3.6+532=60/3.6 t

t=33 s.